OpenStudy (anonymous):

Estimate the value of the following convergent series with an absolute error less than 10^-3.

OpenStudy (anonymous):

\[\sum_{k=1}^{\infty} \frac{ (-1)^{k} }{ k ^{k} }\]

OpenStudy (anonymous):

I can't tell what \[a _{k}\] is supposed to be..\[\frac{ 1 }{ k ^{k} }\] ?

OpenStudy (anonymous):

okay, i can rewrite that as \[\frac{ (-1)^{k} }{ k ^{k} } as (-1)\left( \frac{ 1 }{ k } \right)^{k}\] ?right?

OpenStudy (anonymous):

help please?